OpenStudy (anonymous):

int from -1to1 1/x dx

OpenStudy (anonymous):

is that a question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is the solution?

OpenStudy (anonymous):

let u=1-x du=-du int(1-u)/udu =-int1/u-1du =1-x-ln(1-x)+c

OpenStudy (anonymous):

correct

OpenStudy (turingtest):

let's not let's instead note that 1/x is an odd function

OpenStudy (turingtest):

\[\int_{-a}^{a}f(x)dx=0\]if f(x) is odd

OpenStudy (anonymous):

@TuringTest if f(x) is odd then what?

OpenStudy (turingtest):

then the above theorem holds

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@TuringTest can you please show me the steps of this question?

OpenStudy (anonymous):

wow u dont give me a meda

OpenStudy (turingtest):

there is a theorem that\[\text{ iff }f(x)\text{ is odd, then }\int_{-a}^{a}f(x)dx=0\]do you know what the definition of an odd function is @woundedtiger40 ?

OpenStudy (anonymous):

I never heard of it :(

OpenStudy (turingtest):

a function f(x) is odd if\[f(-x)=f(x)\] so functions like x, x^3, and 1/x are odd, because\[f(-x)=\frac1{-x}=-\frac1x=-f(x)\]

OpenStudy (anonymous):

but when I try to solve this question in maple it says undefined

OpenStudy (turingtest):

because you have a singularity at x=0, oops forgot about that :/

OpenStudy (anonymous):

I have found something about it at http://planetmath.org/encyclopedia/IntegralOfOddFunction.html but what exactly is a an odd function?

OpenStudy (anonymous):

gotcha

OpenStudy (turingtest):

I just defined an odd function above when f(-x)=-f(x) the function is odd plot a few odd functions on wolfram like x, x^3, 1/x, etc into wolfram and you will see they have a certain symmetry

OpenStudy (anonymous):

thanks dude

OpenStudy (turingtest):

but because we have a singularity at x=0 this is an improper integral, so we need to convert this to\[\lim_{a\to0^-}\int_{-1}^{a}\frac{dx}x+\lim_{a\to0^+}\int_{a}^{1}\frac{dx}x\]in other words, my trick won't work here because 1/x is undefined at x=0 do yuo know anything about improper integrals yet?

OpenStudy (turingtest):

so actually, this integral diverges unless you use the Cauchy-Principle Value, which I will not explain in detail, but basically allows you to use my odd function trick in this case i.e.\[\int_{-1}^{1}\frac{dx}x\text{ diverges, but }PV\int_{-1}^{1}\frac{dx}x=0\]